Talia’s Tutus bought a new sewing machine for $85,000 that will be depreciated over 5 years using double-declining-balance depreciation with a switch to straight-line.

Required:

a. Find the depreciation charge each year.

b. If the sewing machine is sold after 3 years for $35,000, what will be the after-tax proceeds on the sale if the firm’s tax bracket is 35%?

Part A:

YEar Depreciation
1
2
3
4
5

Part B:

After-Tax Proceeds: ______________________
Answer #1

a)

Double declining rate=2*1/usefule life of asset

=2*1/5

=40%

b)

Sale value =35,000

(-)Book value at the end of 3rd year=18,360

Profit =16,640

Less:tax@35% =5,824

After tax proceeds =35,000-5,824

=$29,176